OpenStudy (anonymous):

if 36 squared =1296 and 15 squared = 225 what is (36+15)(36-15)

OpenStudy (anonymous):

a^2 - b^2 = (a + b)(a - b) Which is an important rule to remember! So, substituting a = 36 and b = 15 (36 + 15)(36 - 15) = 36^2 - 15^2 = 1296 - 225 = 1071